OpenStudy (anonymous):

Which of the following points is on the graph of the equation y = x2 ? A. (-2, 4) B. (-5, 10) C. (5, 10) D. (4, 8)

OpenStudy (anonymous):

y=x^2 when we put x=-2 in given eqn, y=(-2)^2 y=4.

OpenStudy (anonymous):

I agree...it is (-2,4)

OpenStudy (mathstudent55):

Try each ordered pair in the equation. The one that makes the equation true is the answer.